# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v2

@@ Abstract
Symmetry is one of the most significant foundational principles underlying nature. The resource theory of asymmetry (RTA) is a resource-theoretic framework for investigating asymmetry as a resource to break constraints imposed by symmetries. It has recently undergone significant developments, resulting in applications in a variety of research areas since symmetry and its breaking are ubiquitous in physics. Nevertheless, the resource conversion theory at the core of RTA remains incomplete. In the independent and identically distributed (i.i.d.) setup, where identical copies of a state are converted to identical copies of another state, conversion theory among pure states has been completed only for $U(1)$ group and finite groups. Here, we establish an i.i.d. conversion theory among any pure states in RTA for any continuous symmetry described by a compact Lie group, which includes the cases where multiple conserved quantities are involved. We show that the quantum geometric tensor is an asymmetry monotone for pure states that determines the optimal approximate asymptotic conversion rate. Our formulation achieves a unified understanding of conversion rates in prior studies for different symmetries. As a corollary of the formula, we also affirmatively prove the Marvian-Spekkens conjecture on reversible asymptotic convertibility in RTA, which has remained unproven for a decade.
@@ I Introduction
Symmetry lies at the core of our fundamental understanding of nature. It governs conservation laws and interactions, influencing diverse areas of modern physics. A new perspective on symmetry has been recently developed [1, 2, 3, 4, 5, 6] in the context of quantum resource theories [7]. Quantum resource theory is a versatile framework for investigating a situation where freely implementable operations and preparable states are restricted. These limitations can often be circumvented with the help of states that cannot be freely prepared, which are regarded as resources. Building on the pioneering development of entanglement theory [8], the framework of quantum resource theories expanded to broader concepts, such as coherence [9, 10, 11], athermality [12, 13], and asymmetry [1, 2, 3, 4, 5, 6]. The resource theory of asymmetry (RTA) [1, 2, 3, 4, 5, 6], which was also called the resource theory of quantum reference frame in earlier studies [1, 2, 3], is a branch of quantum resource theories where the restriction is imposed by a symmetry group, and asymmetry that breaks the symmetry is treated as a resource.
Symmetry and its breaking are ubiquitous in nature, playing crucial roles in numerous physical phenomena. Consequently, studies in RTA have highlighted its applications in broad research topics, such as super-selection rules [1, 2, 3], acceleration of quantum operations [14], no-broadcasting theorems for quantum coherence [15, 16], quantum thermodynamics [17, 18, 19, 20], creation of accurate clocks [19, 21], quantum error correction [22, 23, 24, 25, 26, 27, 28], gate implementation in quantum computation [29, 30, 25, 27], quantum measurements [31, 32, 5, 33, 27], and black hole evapolation [25, 27].
Despite the significant progress in RTA, its resource manipulation theory, a central issue in any resource theory, is still incomplete. In the independent and identically distributed (i.i.d.) setup where identical copies of a state are converted to identical copies of another state, the exact and approximate settings are investigated in RTA, where exact means no error is allowed in conversion, while approximate implies asymptotically vanishing error is allowed in conversion. For a discrete symmetry described by a finite group, a recent study [34] established both exact and approximate i.i.d. conversion theory for pure states, which generalizes the result on $\mathbb{Z}_{2}$ in [2]. For a continuous symmetry described by a Lie group, however, little is known; an approximate conversion theory is completed only for $U(1)$ in [19, 21] generalizing the prior results [35, 36, 2], and partial results on specific examples on restricted sets of pure states are obtained for other groups [2, 37]. This means that a comprehensive theory is still lacking for interesting cases involving multiple conserved quantities where non-commutativity is essential. Even in a special case where the asymptotic conversion is reversible, a definitive understanding remains elusive. A necessary and sufficient condition for conversion to be asymptotically reversible was conjectured by Marvian and Spekkens in [6, 4]. However, this Marvian-Spekkens conjecture had remained unproven for a decade. It is worth mentioning that the asymptotic conversion rate in RTA cannot be derived from the approach in general resource theory for reversible conversion [38] using the generalized Stein’s lemma [39, 40, 41, 42]. This is because this framework does not work well for RTA since the regularized relative entropy vanishes in RTA [3], as emphasized in [38, 41]. As such, a unified understanding of asymptotic conversion theory has not yet been established.
Establishing an asymptotic resource conversion theory is essentially inseparable from identifying an important resource measure. Entanglement entropy in entanglement theory and the Helmholtz free energy in quantum thermodynamics are of distinguished significance as they characterize the asymptotic conversion rates of respective resources [43, 12]. In RTA, various asymmetry measures are proposed, such as $G$-asymmetry [44], the relative entropy of $G$-frameness [3], purity of coherence [19], the quantum Fisher information [21], the quantum Fisher information matrix [45, 46], and the metric adjusted skew information [47, 48, 49] and its extension [50]. Among others, the quantum Fisher information plays a significant role in RTA for $U(1)$ group since the asymptotic conversion rate among pure states is determined by the ratio of the quantum Fisher information [21]. However, it remains elusive how to extend this result to other continuous symmetries.
In this paper, we establish the i.i.d. approximate asymptotic conversion theory among pure states in RTA for an arbitrary compact Lie group. The asymptotic conversion rate is shown to be determined by the quantum geometric tensor (QGT), a local geometric characteristic of rays in a Hilbert space. The QGT is a complex Hermitian tensor whose real symmetric part corresponds to the quantum Fisher information matrix, while the imaginary anti-symmetric part gives the Berry curvature [51, 52]. For a Lie group $G$ with dimension $\dim G$, the QGT is a $\dim G\times\dim G$ Hermitian matrix, which is equal to a non-symmetrized covariance matrix for generators of symmetry transformations. Our formula shows that the asymptotic conversion rate is determined by a finite number of matrix inequalities among the QGTs, where each inequality corresponds to a connected component of a compact Lie group.
The proof of our formula on conversion rate consists of the converse and direct parts, i.e., the impossibility of achieving the conversion rate beyond the formula and the achievability of the optimal conversion rate. The converse part is a consequence of the fact that QGT and its asymptotic rate are valid asymmetry measures for pure states, proven by using the properties of Petz’s monotone metrics [53] and a technique analyzing its asymptotic behavior [50]. In the proof of the direct part, we adopt results on quantum local asymptotic normality (QLAN) [54, 55, 56, 57] to extract the asymptotic characteristic of i.i.d. copies of a state, and construct quantum channels that asymptotically transform original states into the desired states with exponentially small failure probability. Our result shows that the QGT plays a distinguished role in RTA for an arbitrary compact Lie group that determines the asymptotic conversion rate.
Our formula is valid for any continuous symmetry described by a compact Lie group. To illustrate its wide applicability, we provide four examples. First, since the asymptotic conversion theory among pure states has been completed in prior research exclusively for the case of $U(1)$ [21], we confirm that this result immediately follows from our formula. Second, we provide a necessary and sufficient condition for conversion to be reversible. As a consequence, we affirmatively prove the Marvian-Spekkens conjecture, which had been unsolved for the past decade until today. Third, we analyze examples for $U(1)$ and $SU(2)$ in which the asymptotic conversion is irreversible. The conversion rate for the example in $SU(2)$ is shown to be consistent with the prior study on a specific class of pure states in [2]. Fourth, we calculate the conversion rate for finite groups. Although a finite group is not a Lie group  [^1] Formally, a finite group can be regarded as a Lie group since it can be viewed as a zero-dimensional smooth manifold. However, in our terminology, we consider Lie groups to have a dimension greater than zero., it is shown that our formula is applicable by appending a trivial $U(1)$ phase. We find a divergence in conversion rate, reproducing the prior result on the approximate conversion rate in RTA for finite groups in [34]. Based on our proof of the direct part, we also provide an intuitive explanation of the reason why the approximate conversion rate can diverge in RTA for finite groups.
Our analysis reveals that, for the conversion rate to be nonzero, there must be an inclusion relationship between the symmetry subgroups of the initial and final states. Here, a symmetry subgroup is defined as the set of group transformations that leave a state invariant, serving as a generalization of the period of a state in the $U(1)$ case. However, we show that this stringent constraint imposed by these symmetry subgroups can be circumvented in an alternative scenario where a sublinear number of additional states is available as an auxiliary resource.
Our main result for the conversion rate fully establishes an i.i.d. conversion theory for pure states. We also attempt to extend the analysis to mixed states. Specifically, we study distillable asymmetry and asymmetry cost, which are defined as the optimal rates for converting a mixed state to and from a given pure reference state, respectively. Central to our analysis of distillable asymmetry is a newly introduced asymmetry measure that extends QGT to mixed states. By using the monotonicity of this measure, we derive an upper bound on distillable asymmetry, which generalizes the result in an earlier study [19] in the $U(1)$ case. We also obtain an upper bound on the asymmetry cost and demonstrate that a known result [21] for the $U(1)$ case cannot be directly generalized to arbitrary compact Lie groups.
@@
The rest of this paper is organized as follows: We first review the framework of RTA and the definition of the QGT in Sec. II. The main result of this paper, i.e., the approximate asymptotic conversion rate among pure states in RTA for an arbitrary Lie group, is presented in Sec. III. We also apply our formula for RTA for several setups, providing a unified understanding of prior studies on conversion rates in RTA for different groups [2, 4, 6, 37, 21, 34], and proves the Marvian-Spekkens conjecture [6, 4]. In Sec. IV, we show that the QGT is asymmetry monotone for pure states in one-shot and asymptotic conversion. As a consequence, we prove an inequality providing a bound for the asymptotic conversion rate, i.e., the converse part of the main theorem. In Sec. V, we show the optimality of this bound, i.e., the direct part of the main theorem. In Sec. VI, we prove that a sublinear number of additional states can improve the conversion rate by eliminating the restriction due to symmetry subgroups. In Sec. VII, we investigate mixed-state asymmetry and prove upper bounds on distillable asymmetry and asymmetry cost.
@@ II.1 Resource theory of asymmetry
As a simple yet important characterization of convertibility among states, the symmetry subgroup is known, which is the set of group elements that leave the state invariant, i.e., $\mathrm{Sym}_{G}(\rho)\coloneqq\{g\in G\mid U(g)\rho U(g)^{\dagger}=\rho\}$ for the input state $\rho$, and similarly for the output state $\sigma$. It is shown in [4] that if a state $\rho$ can be convertible to a state $\sigma$ via some $G$-covariant operation without error, the symmetry subgroups must satisfy $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\sigma)$.
@@ II.2 Quantum geometric tensor
Here, we introduce the QGT [51, 52]. For a given parametric family of pure states |$\ket{\xi(\bm{\lambda})}$ with $\bm{\lambda}\in\mathbb{R}^{m}$, QGT at |$\ket{\psi}\coloneqq\ket{\xi(\bm{0})}$ is an $m\times m$ Hermitian matrix $\mathcal{Q}^{\psi,\xi}$ whose matrix elements are defined by
[…]
where |$\ket{\partial_{\mu}\psi}\coloneqq\frac{\partial}{\partial\lambda^{\mu}}\ket{\xi(\bm{\lambda})}|_{\bm{\lambda}=\bm{0}}$. The real and symmetric part of QGT is the quantum Fisher information matrix, while the imaginary anti-symmetric part corresponds to the Berry curvature [51, 52, 61, 62].
In RTA for a Lie group, we can define QGT using a natural parametric family of a pure state defined with a projective unitary representation. Let $G$ be a Lie group, and $\dim G$ denote its dimension as a smooth manifold. Elements in the neighborhood of the identity $e\in G$ can be parametrized as $g(\bm{\lambda})=e^{\mathrm{i}\sum_{\mu=1}^{\dim G}\lambda^{\mu}A_{\mu}}$ by using a basis $\{A_{\mu}\}_{\mu=1}^{\dim G}$ of the Lie algebra $\mathfrak{g}$, where the convention in physics for the definition of Lie algebra is adopted, which differs from that in mathematics by a factor of the imaginary unit. Given a projective unitary representation $U$, we define |$\ket{\xi(\bm{\lambda})}\coloneqq U(g(\bm{\lambda}))\ket{\psi}$ for a pure state |$\ket{\psi}$. For simplicity, we assume that the map $U$ is differentiable, which follows from continuity in the case of unitary representations. We treat the cases where $U$ is continuous but not differentiable in Appendix A.1. For a given $U$, we introduce Hermitian operators
for $\mu=1,\cdots,\dim G$, which corresponds to $L(A_{\mu})$, where $L$ is the Lie algebra representation defined by $L(A)\coloneqq-\mathrm{i}\left.\frac{\mathrm{d}}{\mathrm{d}t}U(e^{\mathrm{i}tA})\right|_{t=0}$. Then the QGT for |$\ket{\xi(\bm{\lambda})}$ at |$\ket{\psi}$, which we simply denote by $\mathcal{Q}^{\psi}$, is equal to the non-symmetrized covariance matrix for $\bm{X}\coloneqq\{X_{\mu}\}_{\mu=1}^{\dim G}$:
@@ III Main results
Let $U$ and $U^{\prime}$ be projective unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. The conversion rate from a pure state |$\ket{\psi}\in\mathcal{H}$ to another pure state |ϕketitalic-$\ket{\phi}\in\mathcal{H}^{\prime}$ is calculated as follows: (I) When $\mathrm{Sym}_{G}(\psi)\not\subset\mathrm{Sym}_{G}(\phi)$, $R(\psi\to\phi)=0$. (II) When $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$,

### v3

@@ Abstract
Quantifying physical concepts in terms of the ultimate performance of a given task has been central to theoretical progress, as illustrated by thermodynamic entropy and entanglement entropy, which respectively quantify irreversibility and quantum correlations. Symmetry breaking is equally universal, yet lacks such an operational quantification. While an operational characterization of symmetry breaking through asymptotic state-conversion efficiency is a central goal of the resource theory of asymmetry (RTA), such a characterization has so far been completed only for the $U(1)$ group among continuous symmetries. Here, we identify the complete measure of symmetry breaking for a general continuous symmetry described by any compact Lie group. Specifically, we show that the asymptotic conversion rate between many copies of pure states in RTA is determined by the quantum geometric tensor, thereby establishing it as the complete measure of symmetry breaking. As an immediate consequence of our conversion rate formula, we also resolve the Marvian-Spekkens conjecture on conditions for reversible conversion in RTA, which has remained unproven for over a decade. By applying our analysis to a standard setup in quantum thermodynamics, we show that asymptotic state conversion under thermal operations generally requires macroscopic coherence in the thermodynamic limit.
@@ I Introduction
Quantifying fundamental concepts through the ultimate performance of a given task has driven theoretical advances in physics. Thermodynamic entropy, emerging from the study of the ultimate limits of state transformation under adiabatic operations, serves as a measure of irreversibility and has led to a modern formulation of the second law of thermodynamics [1]. Likewise, entanglement entropy was discovered as a measure of quantum correlations that fully characterizes the constraints on state conversion imposed by local operations and classical communication (LOCC) [2]. Its relevance now extends beyond quantum information science, playing a pivotal role in developments across various fields of physics, including condensed matter physics [3, 4] and high-energy physics [5].
However, symmetry and its breaking, equally universal to irreversibility and quantum correlations, still lack such an operational characterization. They are cornerstones of physics, which play a vital role in characterizing natural phenomena across almost every modern field. Considering its ubiquity, quantifying symmetry breaking in the presence of symmetry in terms of operational limits will be a foundation to yield profound theoretical insights and a far-reaching impact, paralleling the influential role of entanglement entropy in modern physics.
The operational quantification of fundamental concepts constitutes a central challenge in resource theories [6], which serve as versatile frameworks generalizing entanglement theory and thermodynamics. Their major goal is to identify the quantity that fully specifies the optimal asymptotic conversion rate between many identical copies of quantum states by given allowed operations, which is called a complete measure [7, 8]. Entanglement entropy, von Neumann entropy, and Helmholtz free energy fulfill this role for LOCC operations [2], adiabatic operations [9], and isothermal operations [10], respectively. Similarly, identifying an analogous measure for symmetry breaking necessitates the development of a suitable resource theory.
The resource theory of asymmetry (RTA) [11, 12, 13, 14, 15, 16], previously referred to as the resource theory of quantum reference frames [11, 12, 13], provides a rigorous framework for quantifying symmetry breaking. However, despite active research in RTA for over a decade [11, 12, 13, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 28, 30, 34, 35, 15, 36, 30, 14, 16, 37], progress toward identifying a complete measure for symmetry breaking remains limited. In particular, for continuous symmetries, a complete measure has been established only for the simplest $U(1)$ group [12, 23, 24], while nothing is currently known about any other continuous groups beyond a few highly specific examples of states [12, 38]. Consequently, while the RTA for the $U(1)$ group has led to applications in a wide range of fields, including quantum thermodynamics [20, 21, 23, 22], measurements [34, 35, 15, 36, 30, 39], quantum computing [32, 33, 28, 30], error-correcting codes [25, 26, 27, 28, 29, 30, 31], and black hole physics [28, 30], it has a limited impact on research domains dictated by non-Abelian Lie groups, such as $SU(2)$ spin‑rotation symmetry in quantum magnets [40, 41] and cold‑atom Hubbard models [42], and the $SU(4)$ symmetry realized in multicomponent fractional quantum Hall systems in graphene [43]. Therefore, identifying fundamental measures for symmetry breaking for non-Abelian Lie groups is an urgent task to open new avenues in these areas.
In this paper, we identify the complete measure of symmetry breaking for general continuous symmetries described by any compact Lie groups. Specifically, we show that the quantum geometric tensor (QGT) [44, 45], consisting of the quantum Fisher information matrix and the Berry curvature, fully determines the asymptotic conversion rate between independent and identically distributed (i.i.d.) pure states in RTA. Interestingly, the QGT has been studied in a different context of topological phases of matter [46, 47, 48, 49]. Our result reveals its operational interpretation by establishing it as the complete measure of symmetry breaking in RTA, and thus provides information-theoretic foundations for the use of QGT, much like entanglement entropy in entanglement theory.
Our result also provides a fundamental formula for the exploitation of symmetry breaking as a resource in quantum technologies. Concretely, it presents the ultimate limits of distillation [50], i.e., converting many low-quality resource states into a few high-quality ones, which has played a critical role in the field of entanglement, as illustrated in applications such as quantum cryptography [51], distributed quantum computing [52], and quantum internet [53]. Moreover, since our formula is expressed through a finite set of matrix inequalities, it yields a computable prediction of the fundamental efficiency bound of the distillation processes. As an immediate consequence of the formula, we also derive a necessary and sufficient condition for reversible conversion, which resolves a long-standing open problem known as the Marvian-Spekkens conjecture [14, 16].
The development of tools for quantifying symmetry breaking in RTA enables applications across diverse scenarios where dynamics are constrained by conservation laws. A particularly important case is energy conservation, which is equivalently expressed as time-translation symmetry. The breaking of this symmetry—namely, coherent superpositions of energy eigenstates with different eigenvalues—represents a fundamental resource, commonly referred to as energy coherence. Quantum thermodynamics [10, 20, 54, 55, 56, 57, 58] offers a natural setting for studying this resource, as its central objective is to determine which state transformations are achievable under thermal operations—evolutions driven by interactions with a heat bath that preserve total energy. For states without any energy coherence, called quasiclassical states, the non-equilibrium free energy fully characterizes state convertibility [10]. However, the situation for general, non-quasiclassical states remains unsettled [57, 58]. The difficulty arises from the fact that thermal operations cannot generate energy coherence, which must therefore be regarded as an independent resource, distinct from non-equilibrium free energy.
Existing approaches [10, 57, 58] have often quantified the required coherence in terms of the energy range of an external system that provides energy coherence, concluding that only a small source of coherence is needed in the thermodynamic limit. However, this perspective does not provide a refined characterization of state-specific resource requirements, as it reflects only the specifications of the external system. Building on the establishment of the QGT as a complete measure of asymmetry for a broad class of continuous symmetries, we revisit this problem. We demonstrate that, contrary to the conventional view, certain state conversions via thermal operations require a macroscopic amount of asymmetry—including energy coherence—that scales extensively with system size when quantified with the QGT. This finding highlights that the choice of asymmetry measure is a decisive factor for advancing our understanding of resource requirements in thermodynamic processes.
The significance of quantifying symmetry breaking is not limited to the study of the dynamics of symmetric systems, but it also extends to the investigation of quantum reference frames. Any physical operation is fundamentally defined relative to a reference structure, such as a time origin or a Cartesian frame. If such a structure is absent or misaligned, operations cannot, in general, be perfectly implemented. This effective lack of reference gives rise to superselection rules [11], which restrict the set of feasible operations. A symmetry-breaking quantum state therefore serves as a valuable resource for overcoming these restrictions. For example, a state that is not invariant under time translation can serve as a quantum clock, specifying the time origin. Similarly, a state that is not invariant under spatial rotations can function as a quantum gyroscope, defining the orientation of a Cartesian frame. Building on this perspective, the RTA has in fact been developed as a theory of quantum reference frames in earlier works [11, 12, 13].
When symmetry breaking is regarded as a resource for correcting misalignment, the ability to manipulate such resources becomes essential. In particular, efficient communication requires systematic conversions of these resources. For instance, when quantum communication channels are limited, it is essential to convert existing resources into high-quality states, thereby maximizing the reference-frame information transmitted per channel use. Conversely, if the supply of symmetry-breaking resources is restricted but the goal is to distribute them among many parties, the priority shifts to producing as many resource states as possible, even at the expense of their individual quality. The conversion theory within the RTA directly applies to such scenarios, providing explicit protocols for optimal conversions. As a demonstration, we show that the conversion theory enables a systematic definition of standardized reference states for continuous symmetries associated with semisimple compact Lie groups.
@@
The rest of this paper is organized as follows: We first review the framework of RTA and the definition of the QGT in Sec. II. The main result of this paper, i.e., the approximate asymptotic conversion rate among pure states in RTA for an arbitrary Lie group expressed in terms of QGTs, is presented in Sec. III. We also apply our formula for RTA for several setups, providing a unified understanding of prior studies on conversion rates in RTA for different groups [12, 14, 16, 38, 24, 59], and proves the Marvian-Spekkens conjecture [16, 14]. In Sec. IV, we show that the QGT is a resource measure in the RTA for pure states under one-shot and asymptotic conversion. As a consequence, we prove an inequality providing a bound for the conversion rate. In Sec. V, we show the optimality of this bound, which completes the proof of the main result. In Sec. VI, we show that a sublinear number of additional resource states can significantly improve the conversion rate. In Sec. VII, we extend our analysis to mixed states to obtain the limitations of dilution and distillation processes. In Sec. VIII, as an application of our results, we investigate resource requirements in quantum thermodynamics with multiple conserved quantities, showing that certain state conversions require a macroscopic amount of asymmetry, including coherence, in the thermodynamic limit. In Sec. IX, we review the connection between symmetry breaking and quantum reference frames, and show that our main result enables us to introduce a standardized reference state of reference frames.
@@ II.1 Resource theory of asymmetry
As a simple yet important characterization of convertibility among states, the symmetry subgroup is known, which is the set of group elements that leave the state invariant, i.e.,
$\mathrm{Sym}_{G}(\rho)\coloneqq\{g\in G\mid U(g)\rho U(g)^{\dagger}=\rho\}$ (1)
for the input state $\rho$, and similarly for the output state $\sigma$. It is shown in [14] that if a state $\rho$ can be convertible to a state $\sigma$ via some $G$-covariant operation without error, the symmetry subgroups must satisfy $\mathrm{Sym}_{G}(\rho)\subset\mathrm{Sym}_{G}(\sigma)$.
@@ III Main results
[…]
@@
Interestingly, the conversion rate formula can also be expressed by using the quantum max-relative entropy. For positive semi-definite operators $\rho$ and $\sigma$, the quantum max-relative entropy [66, 67] of $\rho$ with respect to $\sigma$ is defined by
$D_{\max}(\rho\|\sigma)\coloneqq\inf\{\lambda\in\mathbb{R}\mid\rho\leq 2^{\lambda}\sigma\}.$ (10)
Accordingly, the formula for a compact Lie group in Eq. (7) can be rewritten as
$R(\psi\to\phi)=2^{-\max_{i}D_{\max}\left(\mathcal{Q}^{\phi_{g_{i}}}\middle\|\mathcal{Q}^{\psi_{g_{i}}}\right)},$ (11)
where we have introduced the shorthand notations: $\psi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}(\psi)$ and $\phi_{g_{i}}\coloneqq\mathcal{U}_{g_{i}}^{\prime}(\phi)$. In particular, for a connected compact Lie group, this expression is further simplified to
$R(\psi\to\phi)=2^{-D_{\max}(\mathcal{Q}^{\phi}\|\mathcal{Q}^{\psi})}.$ (12)
It is worth emphasizing that, despite the widespread use of the quantum max-relative entropy in various contexts—including entanglement theory [66, 68], quantum asymptotic equipartition property [69], quantum cryptography [70], entropic uncertainty relations [71] and quantum coherence [72]—it has been used exclusively as a divergence measure between (sub)normalized states in these applications. In contrast, to the best of our knowledge, this study is the first to apply the max-relative entropy between resource measures themselves and to connect it directly to the operationally meaningful conversion rate in the RTA. In this sense, our study extends its significance beyond established uses.